OpenStudy (anonymous):

Janet's gross pay is $750, from which $31 is deducted for OASDI, $10 for Medicare, and $14 for income tax. What is her net pay?

OpenStudy (anonymous):

Net pay = 750 -(31+10+14) = 750 -55= 695 dollars @TylerBunner